Tech Staff Summary, 1/4/2000

Current Priorities

Y2k
No problems were encountered.

GPS
Obion HS is dead; need a new site
Piggot has gone out for bid
Ripley will be nearly impossible without sacrificing perfect sky-view
Chris will show a potential site at Ripley with possibly acceptable sky view to one of the PI's.
Chris will scout around north of Pickwick for some potential sites
Chris will scout around north of New Madrid for some potential sites

Digital Backbone
Arch has sent the Water District Contract to the UM lawyers. The lawyers think an acceptable rewording of the hold-harmless clause can be worked out within a month. If so, this aggreement will be used as a model for others that will be needed. If not, this project is more or less dead.
Greg will order two sets of radios and dishes for both analog and digital testing.
Chris and Jim will verify that the Arp water tower is actually on state property since the new markers appear as if it isn't.
Greg will provide Mitch with lat/lon pairs of sites for the preferred path and for sites in one or two options.
Mitch will get specs for routers to Greg.
If the Memphis Library Tower in Arkansas will work, we may be able to avoid putting up a tower at Shelby Forest.

Marked Tree Broadbands
Jim will replace HBAR with the MD-3 sensor so that we can fix the strange long period signal on the HBAR north channel.
Mitch has been in contact with the tech at Gosnell HS near Blytheville to get approval for this as the MD-3 site. The tech is taking it to the board and we are awaiting the results.

ET
John is moving rapidly on getting permissions for the northern node sites. He and Mitch have been trading letters with various DOI agencies. John will get new telemetry paths to Greg as soon as we know what they are.
Mitch's trip to Morristown got cancelled. He will reschedule for sometime in late January or early February such that Greg can go too.
The water leak for the S-13 Horizontals remains a mystery. Repairs to BCRT.N, BHT.N, GMG.E, and RCG.E have been delayed until this can be solved.
We need to make sure that John in not slowed down by lack of support and hardware on our end. Greg will get John to show Jim and Hoyt how to calibrate sensors. Jim will make sure other hardware (e.g. discriminators, receivers, etc) is being worked on.
John is awaiting permission to move RICH.

CERI power systems
Greg has cleaned out the racks. Mitch will provide a sketch on what computers go where in the racks.
Mitch/Steve will get the appropriate cable extensions.

Second tier priorities:

Autozone
Dorman has $2800 to put toward this; enough for the pc, digitizer, and mux.
Mitch will submit a budget and acquire the hardware.
Greg will refurbish one of the old GPS clocks that supplies the wrong time code. While this time code won't work with the DAQ system, we will record it on one channel in the event absolute time is needed. Data will be stamped with the pc time since, under normal circumstances, only relative time is important.
Jim will service the free-field sensor. Then, after the DAQ is up and we get some data, he will service the other sensors as required.

Other Broadbands
Greg has repaired and tested the CMG-3/DM-24 packages and await testing in Mitch's office.
Potential sites for these sensors are Sewanee and Morristown, TN.
Where the hell is the borehole sensor?
Status of remaining (3) CMG40-T's slated for ET?

Aftershock Array
We are unaware of anything that needs doing here. Presumably Bodin has it under control.

MEQ's
The L-4's from NC will replace the 4.5Hz geophones currently being used.

ISIS upgrade
As time permits, stations will be upgraded to the new 4-channel system with emphasis on the backbone stations (this'll make Harley happy to get clipped data).
New Madrid node still needs the new electronics.

Mitch will attempt to find a home for the geophysical prospecting equipment

As time permits priorities.

Panda/PandaII
No deployments are firm. Proposals have been submitted which could result in deployments no sooner than 10/2000. PII on south campus should be inventoried. Then, the 15 PII packages at CERI should be fixed. Then, we can worry about PI.

SFTN power upgrade
The glitches don't bother Paul so they don't bother us. Greg spent a day and a half making the upgrade. It will take about 30 minutes to install. When paul is ready to tag along, we'll install it.

SFTN general upgrade
This would be very nice, but makes Mitch nervous. It requires pulling the sensor.

Analog backbone will be reconfigured simultaneously with digital backbone installation.

EBZ/PWLA dropouts have apparently been fixed.

LNXT node
Secure propane tank.
Secure racks.
replace coax jumpers
replace rg-58 unshielded jumpers

MKTA node

NMAD node

Station book. Jim has good records for each station but it would be nice to have an IRIS style station book.

Backburner priorities

Status channels. Greg has a scheme, but the usefulness is dubious without a digital backbone. Mitch still has to write the decoding software (which he won't do without live feeds to test it on).

BB digital telem interference is still a mystery. Continous record of when gaps occurr would be helpful. This would be trivial if there were a digital backbone.

Guralp reset problem is awaiting results from Guralp. One of CERI's DM-16's is in England for this purpose. Wonder if we'll ever see it again?

LNXT node
permit/install L7 and L8

MKTA node
permit/install remaining 3 ISIS

NMAD node